### v6

# Biology-inspired joint distribution neurons based on Hierarchical Correlation Reconstruction for multidirectional propagation of values and densities

Abstract
Recently a million of biological neurons (BNN) has turned out better from modern RL methods in playing pong [1], reminding they are still qualitatively superior e.g. in learning, flexibility and robustness - suggesting to try to improve current artificial e.g. MLP/KAN for better agreement with biological. There is proposed extension of KAN approach to neurons containing model of local joint distribution: $\rho(\mathbf{x})=\sum_{\mathbf{j}\in B}a_{\mathbf{j}}f_{\mathbf{j}}(\mathbf{x})$ for $\mathbf{x}\in[0,1]^{d}$, adding interpretation and information flow control to KAN, and allowing to gradually add missing 3 basic properties of biological: 1) biological axons propagate in both directions [2], while current artificial are focused on unidirectional propagation - joint distribution neurons can repair by substituting some variables, getting conditional values/distributions for the remaining. 2) Animals show risk avoidance [3] requiring to process variance, and generally real world rather needs probabilistic models - the proposed can predict and propagate also distributions as vectors of moments: (expected value, variance) or higher. 3) biological neurons require local training, and beside backpropagation, the proposed allows many additional ways, like direct training, through tensor decomposition, or finally local and very promising: information bottleneck. Proposed approach is very general, can be also used as extension of softmax $\textrm{Pr}\propto\exp(-E)$ e.g. in embeddings of transformer, into their probability distributions working on $(a_{j})$ few moments: $\rho(x)\approx\sum_{j}a_{j}f_{j}(x)$.
Keywords: machine learning, neural networks, Kolmogorov-Arnold Network, joint distribution, conditional distribution, Bayesian Neural Networks, tensor decomposition, mutual information, information bottleneck , HSIC, softmax, embedding
I Introduction
Just connecting a million of biological neurons to electrodes, recently turned out better from modern Reinforcement Learning method in playing pong [1], reminding that biological neurons still have qualitative superiority, especially in learning, flexibility and robustness, summarized further in Fig. 3. There are low level approaches to catch up like spiking neural networks [4], however, beside energy efficiency, without much successes. Biological neurons might be too complicated for practical direct modelling: they work in bursts of spikes, which annihilate when meet, traveling through complex networks in all directions.
Figure 1: Like in MLP/KAN approach, we search for logical neurons supposed to extract essential mathematics hidden in biological neurons - focusing on their usually missing 3 properties: bidirectional propagation [2], also working of distributions e.g. for observed risk avoidance [3], and using local training like information bottleneck [5]. All 3 are natural for neurons containing joint distribution model, practical with HCR approximating joint density as a linear combination: $\rho(\mathbf{x})=\sum_{\mathbf{j}\in B}a_{\mathbf{j}}f_{\mathbf{j}}(\mathbf{x})$ for $a_{\mathbf{j}}$ moments as neuron parameters.
Figure 2: Basic formulas and example for $d=2$ variables HCR neuron, using convenient variable normalization to nearly uniform in $[0,1]$. Neuron contains matrix of moments: $a_{ij}$ (generally order $d$ tensor), allowing to propagate in various directions by substituting some variables and normalizing to get estimated conditional density for the remaining - just permuting indexes to change propagation direction. For value propagation we can take expected value, what restricts prediction to 1-st moment - becoming summation of trained polynomials like in KAN[6], allowing to view HCRNN as extension of KAN e.g. for optional density propagation and change of propagation direction. Additionally, such $a_{ij}$ parameters are actual moments allowing for better interpretation, and inexpensive estimation of mutual information - for training or monitoring of information flow.
Therefore, to catch up with BNNs, it seems more promising to search for higher level MLP[7]/KAN[6]-like logical neurons - trying to extract mathematics hidden in biological neurons, asking if it is e.g. matrix multiplication-like. To understand this mathematics, we focus on 3 fundamental differences, summarized in Fig. 1: uni-directional vs bi-directional propagation of biological axons [2], working only on values vs also on distributions - observed in animals e.g. as risk avoidance [3], and finally BNNs need local training approaches like looking the most promising: information bottleneck.
Neurons containing local joint distribution model would allow to include above features of BNNs, by substituting some variables allows to predict and propagate conditional probabilities - what seems crucial for animals to operate in complex physical world, practice being quite probabilistic. If only in reach at a reasonable cost, evolution should provide organisms ways to predict conditional distributions, preferably also propagating uncertain information through neural networks.
However, modelling joint distributions is generally quite difficult - for practicality needs to be simplified as possible, like representation of joint density as just a linear combination like polynomial, of conveniently normalized variables - we refer to as HCR, leading to simple formulas with e.g. just matrix/tensor products as in Fig. 2, 5 - relatively inexpensive calculations, hence this kind of representation might be hidden in biological neurons.
Figure 3: Summary of differences between artificial (ANN) and biological neural networks (BNN, based on https://www.geeksforgeeks.org/difference-between-ann-and-bnn/) - BNNs are qualitatively superior in terms of learning, flexibility and robustness - just increasing the number of neurons might be insufficient to reach them. To build ANNs closer to BNN capabilities, we should include their neuron-level properties, summarized in Fig. 1, all possible for neurons containing local joint distribution model - allowing to approach complete statistical description of information available to neuron, for organisms build probabilistic models of the world with their networks.
Figure 4: Simple 2/3D examples from HCRNN Wolfram notebook of propagation in any direction based on the shown datasets (points) as conditional expected values, here being degree $m=8$ polynomials.
Figure 5: The proposed HCR neuron and neural network (HCRN, HCRNN) containing local joint distribution model represented in $(a_{\mathbf{j}})_{\mathbf{j}\in B}$ tensor, e.g. $(a_{ijk})_{i,j,k\in\{0,..,m\}}$ for $d=3$ connections. Top: used orthonormal polynomial basis for uniform weight in $[0,1]$, convenient for such normalization to quantiles.
Middle: HCR neuron containing and applying joint distribution model here for $d=3$ variables, and gathered formulas for direct estimation/model update, its application to propagate entire distributions and expected values alone. Such $\rho$ density parametrization can drop below 0, what is usually repaired by calibration e.g. using normalized $\max(\rho,0.1)$ density. However, for neural networks with inter-layer normalization this issue seems negligible, what essentially simplifies calculations to the shown formulas. Propagating only expected values and normalizing, we can use only the marked nominators - as in KAN optimizing nonlinear functions (polynomial here) by including only pairwise dependencies ($a$ with two nonzero indexes), extending to their products to consciously include higher order dependencies.
Bottom: schematic HCR neural network and some training approaches of intermediate layers - which in HCR can be treated as values or their distributions (replacing $f_{i}(u)$ with its $i$-th moment: $\int_{0}^{1}\rho(u)f_{i}(u)du$). There is also visualized tensor decomposition approach - estimate dependencies (e.g. pairwise) for multiple variables and try to automatically decompose it to multiple dependencies of a smaller numbers of variables with algebraic methods.
The HCR acronym stands for Hierarchical Correlation Reconstruction, and already found many applications ([8, 9, 10, 11, 12, 13]) - allowing to hierarchically decompose dependencies into mixed moments $a_{\textbf{j}}$, predict and propagate them - here extended into neural networks as HCRNN  [^1] HCR introduction: https://community.wolfram.com/groups/-/m/t/3017754, HCRNN application: https://community.wolfram.com/groups/-/m/t/3241700 we focus on in this article.
While this approach can be also viewed as just parametrization, allowing universal backpropagation training, it also allows many other training ways - briefly summarized in the bottom of Fig. 5. E.g. having all neuron connections we could directly estimate $a_{\textbf{j}}$ moments between them, or update their current model. Working on joint distributions with easy to estimate mutual information, we can use mentioned local information bottleneck approach - offering many improvements over previously used HSIC [14, 15], like much lower cost and better sensitivity. There is also possible training by tensor decomposition, extending SVD (singular value decomposition) - trying to automatically decompose larger joint distribution into multiple smaller ones.
The discussed approach is very general - allowing e.g. to predict or propagate densities as vector of moments: approximately (expected value, variance) and optionally higher like skewness, kurtosis. It could be used to also modify standard approaches e.g. softmax, embedding-based like transformers: usually they predict single energy-like coefficient $E$ to define probability with softmax: $\textrm{Pr}\propto\exp(-E)$. However, if in embedding it represents e.g. age, single coefficient seems insufficient to describe our incomplete knowledge - in contrast to extension to multiple moment-like coefficients $a_{j}$, HCR representing density as $\rho(x)\approx\sum_{j}f_{j}(x)a_{j}$, their products (calibrated for un-embedding).
This article introduces to HCR from perspective of neural network applications, earlier suggested in [8], to be extended in future e.g. with practical realizations replacing MLP, KAN, or just to extend current architectures e.g. softmax-based approached to work with probability densities.
II HCR neural networks (HCRNN)
This main Section introduces to HCR and discusses it as a basic building block for nextgen neural networks.
Figure 6: 2D example comparison of local basis KDE (kernel density estimation) vs global basis HCR, similar as available with code in HCR Wolfram notebook - trying to model joint density for the dataset points shown on the right. Assuming trivial $\rho=1$ joint density, we would get 0 log-likelihood evaluation (mean $\lg(\rho(x))$ over dataset). Training on a randomly chosen subset and calculating log-likelihood in the remaining subset (cross-validation), we can see local KDE barely exceed this 0 for trivial model (best for $\epsilon\approx 0.03$ kernel width), while global: polynomial and trigonometric easily exceed it - are able to extract crucial generalizing features. The best for $m=4$ polynomials - using popular 4 moments: expected value, variance, skewness and kurtosis. Intuitively, KDE assumes that new points will be close to the old points - what poorly generalizes, in contrast to global features like moments used by HCR.
Figure 7: Top: Visualized part of HCR polynomial $[0,1]$ basis in $d=1$ dimension and $f_{\textbf{j}}(\textbf{x})=\prod_{i=1}^{d}f_{j_{i}}(x_{i})$ product bases for $d=2,3$. E.g. for $d=3$ the assumed joint density becomes $\rho(x,y,z)=\sum_{ijk}a_{ijk}f_{i}(x)f_{j}(y)f_{k}(z)$. As $f_{0}=1$, zero index in $a_{ijk}$ means independence from given variable, hence $a_{000}=1$ corresponds to normalization, $a_{i00},a_{0i0},a_{00i}$ for $i\geq 1$ describe marginal distributions through $i$-th moments. Then $a_{ij0},a_{i0j},a_{0ij}$ for $i,j\geq 1$ describe pairwise dependencies through mixed moments, and finally $a_{ijk}$ for $i,j,k\geq 1$ describe triplewise dependencies. This way we literally get hierarchical correlation reconstruction through moments describing dependencies between increasing numbers of variables, with clear interpretation of coefficients of e.g. trained HCR-based neural network. Bottom: discussed further information theoretic view, allowing inexpensive estimation of e.g. mutual information as sum of square of nontrivial moments between two variables.
II-A Introduction to Hierarchical Correlation Reconstrution
As in copula theory [16], it is convenient to use normalization of variables to nearly uniform distribution in $[0,1]$. It requires transformation through cumulative distribution function (CDF): value $x\to\textrm{CDF}(x)$ becomes its estimated quantile, e.g. $1/2$ for median. This CDF can be modeled with some parametric distribution using parameters estimated from dataset e.g. $x\to$
$\textrm{CDF}_{N(0,1)}((x-E[X])/\sqrt{\textrm{var}(X)})=\textrm{CDF}_{N\left(E[X],\sqrt{\textrm{var}(X)}\right)}(x)$
for Gaussian distribution, or can be empirical distribution function (EDF): $x$ becomes its position in dataset rescaled to $(0,1)$. For neural networks such normalization is usually made in batches [17], here needed to be used between layers (can be skipped in further linearization). In practice should be nearly constant between batches, approximated, parameterized e.g. by Gaussian, put into tables, inversed for backward propagation, etc.
For $d$ normalized variables: $\mathbf{x}\in[0,1]^{d}$, in HCR we represent joint distribution as a linear combination, conveniently in some product basis $B=B^{+}\cup\{\textbf{0}\}$ with $f_{0}(x)=f_{\mathbf{0}}(\mathbf{x})=1$:
$\rho(\mathbf{x})=1+\sum_{\mathbf{j}\in B^{+}}a_{\mathbf{j}}f_{\mathbf{j}}(\mathbf{x})=\sum_{\mathbf{j}\in B}a_{\mathbf{j}}f_{\mathbf{j}}(\mathbf{x})=\sum_{\mathbf{j}\in B}a_{\mathbf{j}}\prod_{i=1}^{d}f_{j_{i}}(x_{i})$ (1)
where $B\{$$^{+}=B\backslash\{\textbf{0}\}$ removes zero corresponding to normalization as $f_{0}=1$, bold fonts denote vectors: $\mathbf{j}=(j_{1},..,j_{d})$.
Assuming orthonormal basis: $\int_{0}^{1}f_{i}(x)f_{j}(x)dx=\delta_{ij}$ static estimation [18] (minimizing mean-squared error between kernel density estimation smoothed sample and discussed parametrization) from $\bar{X}$ dataset becomes just:
$a_{\mathbf{j}}=\frac{1}{|\bar{X}|}\sum_{\mathbf{x}\in\bar{X}}f_{\mathbf{j}}(\mathbf{x})=\frac{1}{|\bar{X}|}\sum_{\mathbf{x}\in\bar{X}}\prod_{i=1}^{d}f_{j_{i}}(x_{i})$ (2)
We assume here orthonormal polynomial basis (rescaled Legendre), allowing to interpret coefficients as moments of normalized variables, being approximately cumulants, starting with:
$f_{0}=1\qquad\textrm{corresponds to normalization}$ (3)
$f_{1}(x)=\sqrt{3}(2x-1)\qquad\sim\textrm{expected value}$
$f_{2}(x)=\sqrt{5}\left(6x^{2}-6x+1\right)\qquad\sim\textrm{variance}$
$f_{3}(x)=\sqrt{7}\left(20x^{3}-30x^{2}+12x-1\right)\qquad\sim\textrm{skewness}$
$f_{4}(x)=3\left(70x^{4}-140x^{3}+90x^{2}-20x+1\right)\qquad\sim\textrm{kurtosis}$
Independent $a_{\mathbf{j}}$ estimation as just average allows to control uncertainty of the found parameters $\propto 1/\sqrt{|\bar{X}|}$.
Alternatively we could use various trigonometric bases (e.g. discrete cosine/sine transform DCT/DST) - especially for periodic variables, localized like B-splines used by the original KAN, from wavelets, finite elements methods. Instead of normalization to uniform in $[0,1]$, we could use a different one e.g. to Gaussian distribution, times Hermite polynomials for orthonormal basis. For discrete variables we can use one-hot encoding, or e.g. its SVD/CCA-based optimization as in [13].
As in Fig. 7 and $f_{0}=1$, $a_{\mathbf{j}}$ coefficients are mixed moments of $\{i:j_{i}\geq 1\}$ variables of nonzero indexes, independent from variables of zero indexes, allowing for literally hierarchical decomposition of statistical dependencies: start with $a_{0..0}=1$ for normalization, add single nonzero index coefficients to describe marginal distributions, then add pairwise dependencies with two nonzero indexes, then triplewise, and so on. For example $a_{2010}$ coefficient would describe dependence between 2nd moment of 1st variable and 1st moment of 3rd variable among $d=4$ variables. Generally the selection of basis $B$ is a difficult question, e.g. to use only pairwise dependencies up to a fixed moment $m$, preferably optimized during training, maybe separately for each neuron or layer. Such decomposition also allows to efficiently work with missing data by using to estimate/update/propagate only $a_{\mathbf{j}}$ coefficients with zero indexes for the missing variables, as $j_{i}=0$ zero index means independence from given variable.
While static estimation averages over dataset with equal weights, for dynamic updating we should increase weights of recent values, e.g. using computationally convenient exponential moving average (EMA) - for some small memory parameter $\eta$:
$a_{\mathbf{j}}\xrightarrow{\mathbf{x}=(x_{1},..,x_{d})}(1-\eta)a_{\mathbf{j}}+\eta\prod_{i}f_{j_{i}}(x_{i})$ (4)
However, modelling (joint) probability density as a linear combination can sometimes lead to negative densities - to avoid this issue, there is usually used calibration: instead of the modelled density $\rho$, use e.g. $\max(\rho,0.1)$ or $\ln(1+\exp(\rho))$ and divide by integral to remain normalized density. However, it makes computations much more complex and costly, especially in higher dimensions - for neural network applications we usually should be able to ignore this issue to simplify calculations, especially working on expected values and normalizing between layers. Therefore, we neglect this issue/calibration in this article, however, it should be remembered, maybe adding calibration for some applications, like final un-embedding into probability distribution e.g. in transformers to find probability distribution e.g. as product over coordinates of calibrated $\rho(x)=\sum_{j}a_{j}f_{j}(x)$.
II-B Conditional distributions and expected value propagation
Having (1) model of joint distribution, to get conditional distribution we need to substitute known variables, and normalize dividing by integral, being 0-th coefficient to make $a_{0}=1$ :
$\rho(x_{1}|x_{2},\ldots,x_{d})=\frac{\sum_{\mathbf{j}}a_{\mathbf{j}}f_{j_{1}}(x_{1})f_{j_{2}}(x_{2})\ldots f_{j_{d}}(x_{d})}{\int_{0}^{1}\sum_{\mathbf{j}}a_{\mathbf{j}}f_{j_{1}}(x_{1})f_{j_{2}}(x_{2})\ldots f_{j_{d}}(x_{d})dx_{1}}=$
$\sum_{j_{1}}f_{j_{1}}(x_{1})\frac{\sum_{j_{2}\ldots j_{d}}a_{j_{1}j_{2}..j_{d}}\,f_{j_{2}}(x_{2})\ldots f_{j_{d}}(x_{d})}{\sum_{j_{2}\ldots j_{d}}a_{0j_{2}..j_{d}}\,f_{j_{2}}(x_{2})\ldots f_{j_{d}}(x_{d})}$ (5)
as $\int_{0}^{1}f_{i}(x)dx=\delta_{i0}$. Such sums for pairwise dependencies use only two nonzero $j_{i}$ indexes (input-output), three for triplewise, and so on. Denominator corresponds to normalization, indeed the fraction becomes 1 for $j_{1}=0$. Examples for $d=2,3$ are shown in Fig. 2, 5 - generally nominator sums over all indexes with the current indexes of predicted variables. Denominator replaces current variables indexes with zeros for normalization, could be removed if having further (inter-layer) normalization.
Here is example of analogous prediction of conditional joint distributions for multiple (2) variables:
$\rho(y,z|x)=\sum_{j_{y}j_{z}}f_{j_{y}}(y)f_{j_{z}}(z)\frac{\sum_{j_{x}}a_{j_{x}j_{y}j_{z}}\,f_{j_{x}}(x)}{\sum_{j_{x}}a_{j_{x}00}\,f_{j_{x}}(x)}$
Working on expected values would remove yzy-z mixed moments, making $E[y,z|x]=E[y|x]E[z|x]$ (can be different for density).
Having such conditional distribution, we can for example calculate expected value e.g. to be propagated by neural networks. For polynomial basis expected values of contributions are: $\int_{0}^{1}xf_{0}(x)dx=1/2,\int_{0}^{1}xf_{1}(x)dx=1/\sqrt{12}$, and zero for higher moments, leading to formulas including only $i=0,1$ normalization and the first moment as in Fig. 5, e.g.:
$E[x|y,z]=\frac{1}{2}+\frac{1}{\sqrt{12}}\frac{\sum_{jk}a_{1jk}f_{j}(y)f_{k}(z)}{\sum_{jk}a_{0jk}f_{j}(y)f_{k}(z)}$ (6)
As further there is rather required CDF normalization which both shifts and rescales, in practice it is sufficient to work on such nominators, marked in Fig. 5, 2.
Restricting it to pairwise dependencies: (single variable of input - single variable of output), similarly to KAN we get summation of trained 1-parameter functions: here polynomials (could be different e.g. B-splines like in KAN) + e.g. approximate fixed CDF for normalization:
${\sum_{jk}a_{1jk}f_{j}(y)f_{k}(z)}\xrightarrow[\textrm{KAN-like}]{\textrm{pairwise only}}\sum_{j}a_{1j0}f_{j}(y)+a_{10j}f_{j}(z)$ (7)
Figure 8: KAN-like example with code from HCRNN Wolfram notebook: trying to find $f(x)=\exp(x_{1}^{2}-x_{2}^{2}-x_{3}^{3}+x_{4}^{4})$ from size 1000 random dataset, using single neuron and direct estimation. We can see it has automatically found the four polynomials hidden behind exponent, however, they are deformed as having values required to fit $[0,1]$ range here.
However, in comparison to KAN, using the proposed HCRNN parametrization we get multiple advantages: it can propagate in any direction (as BNNs), propagate values or probability distributions (as BNNs), interpretation of parameters as mixed moments, consciously add triplewise and higher order dependencies, inexpensive evaluation of modeled mutual information, additional training ways (needed for BNNs), e.g. direct estimation, tensor decomposition, information bottleneck.
Figure 9: Basic HCR-like improvement of polynomial KAN by just replacing standard rescaling normalization with CDF/EDF, what assuming Gaussian distribution means just adding fixed $\textrm{CDF}_{N(0,1)}$ after standard mean subtraction and division by standard deviation. At the top we can see that such quantile normalization leads to more uniform distribution - allowing for better representation with low degree polynomials, which are better at generalization. At the bottom there are results from [19] showing such improvements in cross-validation on MNIST for low degree polynomials. For high degree rescale can be better, but their larger number of parameters is more likely to overfit.
Figure 10: Example of direct prediction of conditional distribution from [20] - here using separate linear regressions for 1st and 2nd moment $a_{j}$ with found coefficients visualized at the bottom, from which we get $\rho(x)=\sum_{j}a_{j}f_{j}(x)$ density as polynomial, which then needs calibration to enforce positive density, and finally normalization can be reverted. We get interpretation by contributions to moments, can evaluate variables by relevance: conditional entropy if predicting from single variable, or novelty: entropy drop if removing single variable from the entire set. While least square linear regression was used, it can be naturally extended to neural networks with MSE estimation of conditional moments, e.g. in the last layer of network originally predicting values, or calculating embeddings e.g. in transformers by a few moments representing its probability distribution.
KAN-like setting as using conditional expected values would include only $(a_{i1})_{i}$ type coefficients - multiple $i$ powers of inputs to predict 1-st moment of output. Reversing propagation direction, it would predict multiple moments as linearly dependent from value. For more sophisticated dependencies and density propagation, we would need to include also $(a_{ij})_{ij}$ dependencies between higher moments.
Figure 8 shows single neuron example learning 4 variable function from sample. Figure 9 shows results from [19] regarding the simplest KAN improvement by just improving normalization from rescaling to quantiles. Fig. 10 shows prediction of conditional distributions from [20] by separate predictions of multiple conditional moments - can be uses as last layer of e.g. KAN, or different neural network e.g. for embedding - as a few conditional moments representing its probability distributions.
II-C Propagation of probability distributions
Let us start with a simple example: that we would like to calculate conditional probability density like previously:
$\rho(x|y)=\sum_{i}f_{i}(x)\frac{\sum_{j}a_{ij}f_{j}(y)}{\sum_{j}a_{0j}f_{j}(y)}$ (8)
but now for $y$ being from $\rho(y)=\sum_{k}b_{k}\,f_{k}(y)$ probability density. So the propagated probability density of $x$ should be $\int_{0}^{1}\rho(x|y)\rho(y)dy$. Approximating with constant denominator, using $\int_{0}^{1}f_{j}(y)f_{k}(y)dy=\delta_{jk}$ and finally normalizing, we get:
$\rho(x)\leftarrow\int_{0}^{1}\rho(x|y)\rho(y)dy\approx\sum_{i}f_{i}(x)\frac{\sum_{j}a_{ij}b_{j}}{\sum_{j}a_{0j}b_{j}}$ (9)
Such constant denominator approximation allows to propagate (in any direction) through HCR neurons not only values, but also entire probability distributions - by just replacing $f_{j}(y)$ for concrete value of $y$, with $b_{j}$ describing its probability distribution. It is natural to generalize, e.g. for $\rho(x|y,z)$ we could replace $f_{j}(y)f_{k}(z)$ with $b_{jk}$ when $\rho(y,z)=\sum_{jk}b_{jk}\,f_{j}(y)f_{k}(z)$:
$\sum_{i}f_{i}(x)\frac{\sum_{j}a_{ijk}\ f_{j}(y)f_{k}(z)}{\sum_{j}a_{0jk}\ f_{j}(y)f_{k}(z)}\quad\mathrel{{\mathop{\leftrightarrows}\limits_{\textrm{density}}^{\textrm{value}}}}\quad\sum_{i}f_{i}(x)\frac{\sum_{j}a_{ijk}\ b_{jk}}{\sum_{j}a_{0jk}\ b_{jk}}$
II-D Tensor decomposition and linearization
Analogously for intermediate layers like in bottom of Fig. 5:
[…]
integrating over the intermediate variables, approximating with constant denominator and normalizing at the end, thanks to basis orthogonality we get Kronecker deltas enforcing equality of intermediate indexes, leading to condition for approximation of higher order tensors with lower order ones, which is generally studied by tensor decomposition field [21] generalizing SVD (singular value decomposition) - hopefully leading to better training approaches (thanks to linear dependencies).
While neural networks require nonlinearity, such tensor approach allows to linearize its intrinsic behavior: calculate nonlinearities in some basis e.g. polynomials only for the outer inputs/outputs $(f_{\mathbf{j}}(\mathbf{x}):\mathbf{j}\in B,\mathbf{x}\in\bar{X})$, including multivariate dependencies. Then treat the entire neural network as a linear transformation of such features (normalization only at both ends - no need for inter-layer), e.g. just changing indexes (like transposition) to modify propagation direction. However, it contains this constant denominator approximation, and it would become a tensor of exponentially increasing size if including all dependencies - it should be combined with reductions like tensor decomposition (into neural network), information bottleneck - working on linear approximations to reduce dimension.
II-E Basis optimization and selection
Another direction is application of the found $a_{\mathbf{j}}$ coefficients, for example to optimize the arbitrarily chosen $\{f_{i}\}$ basis to be able to reduce the number of considered coefficients, also to reduce overfitting issues, e.g. discussed in [13, 22]. For this purpose we can for example treat the current coefficients as a rectangular matrix $M_{j_{1},(j_{2}..j_{d})}:=a_{\mathbf{j}}$ - with blocked all but one indexes for all the considered coefficients in the basis. Now we can use SVD (singular value decomposition): find orthonormal eigenbasis of $MM^{T}=\sum_{i}\sigma_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{T}$ and use $g_{i}=\sum_{j}u_{ij}f_{j}$ as the new basis for one or a few dominant eigenvectors. Similarly we can do for the remaining variables, getting separate or common optimized bases for them.
A more difficult question is basis selection - which $\mathbf{j}\in B$ indexes to use in considered linear combinations for each neuron. Extending all to $m$-th moment/degree for $d$ variables, we would need $(m+1)^{d}=\sum_{k=0}^{d}{d\choose k}m^{k}$ coefficients: 1 for normalization, $dm$ for marginal distributions, $d(d-1)m^{2}/2$ for pairwise, and so on. With proper normalization the coefficients for marginal distributions should be close to 0 - can be neglected. To reduce number of coefficients we can e.g. restrict up to pairwise dependencies ($\approx$ KAN), and/or restrict not individual but summed indexes: $\sum_{i}j_{i}\leq m$ ($\rho$ polynomial degree). Generally we can calculate more coefficients and discard those close to zero in training, maybe use $l^{1}$ normalization to find sparse basis. Using optimized bases as above should allow to reduce $|B|$ size.
II-F Some HCRNN training approaches
A single HCR neuron models multidimensional joint distribution, what is already quite powerful. However, for neural networks the main difficulty is training of the intermediate layers. HCRNN is very flexible also here, below are some approaches:
  • Treat HCRNN as just a parametrization and use standard backpropagation as for MLP, KAN, e.g. with some distance for values, or log-likelihood evaluation for density propagation. It can be mixed with other techniques, e.g. static parameter estimation/update from recent values, online basis optimization and selection, or just optimizing initial parameters to improve further main optimization.
  • Maybe find initial intermediate values by dimensionality reduction like PCA of $\{f_{\mathbf{j}}(x):\mathbf{j}\in B\}$ vectors of features as (nonlinear) products of functions of inputs - further extended into information bottleneck approach.
  • Maybe use propagation in both directions and combine with direct coefficient estimation/update.
  • Maybe use some tensor decomposition techniques - start with estimation of dependencies for a larger set of variables, and use algebraic methods to try to approximate it with multiple lower order tensors.
While backpropagation is available for various parametrizations, such HCRNN with neurons containing joint distribution models bring some additional possibilities - hopefully allowing for faster training, especially with further looking the most promising: information bottleneck approach.
Coefficients of such trained HCRNN remain mixed moments - providing dependency interpretation between input/output and hidden intermediate variables, allowing for multidirectional propagation of values or distributions like in Fig. 13, and its parameters can be further continuously updated e.g. using (4).
III Information bottleneck based training
Let us consider $l$ hidden layer neural network intended to predict $Y$ from $X$ with $\{\theta^{i}\}_{i=0..l}$ sets of parameters:
$X\xrightarrow{\theta^{0}}T^{1}\xrightarrow{\theta^{1}}\ldots\xrightarrow{\theta^{l-1}}T^{l}\xrightarrow{\theta^{l}}\hat{Y}\quad\textrm{predicting }Y$ (10)
While the standard training approach is focused on optimization of neuron parameters $\theta$ used to process the data, alternatively we could try to directly optimize content of $T^{i}$ hidden/intermediate layers, in practice: dataset processed through some of layers (here could be in both directions) - like in image recognition: first layers extract low level features like edges, then some intermediate features, and finally e.g. faces.
Information bottleneck approach ([5, 23]) suggests how to directly optimize the content of intermediate layers. It uses information theory - offering nearly objective evaluation, being invariant to variable permutations and bijections: mutual information I(X;Y)=H(X)+H(Y)H(X,Y)I(X;Y)=H(X)+H(Y)-H(X,Y), being the number of bits (or nits) $X$ on average brings about $Y$, or $Y$ about $X$.
Optimizing intermediate layer of $X\to T\to Y$, obviously it should maximize information about the predicted $Y$: $\max_{T}I(T;Y)$. However, focusing only on prediction would rather lead to overfitting. To prevent that, we can also perform compression: minimize $\min_{T}I(X;T)$ - removing unnecessary information/noise from the input. Finally, information bottleneck approach, for some $\beta>0$, assumes optimization of both:
$\inf_{T}\left(I(X;T)-\beta I(T;Y)\right)$ (11)
III-A Information theory view on HCR
Mutual information has excellent properties: independence to variable permutation and bijections, says how many bits one variable gives on average about the second. However, it is relatively difficult to calculate, requires a joint distribution model - HCR does as $\rho(\mathbf{x})=\sum_{\mathbf{j}\in B}a_{\mathbf{j}}\,f_{\mathbf{j}}(\mathbf{x})$. Using natural logarithm (information in nits) with 1st order approximation $\ln(1+a)\approx a$, and orthogonality of basis, we get simple practical approximation of entropy, formally all differential of normalized variables:
$H(X)=-\int_{[0,1]^{d}}\rho(\mathbf{x})\ln(\rho(\mathbf{x}))d\mathbf{x}\approx-\sum_{\mathbf{j}\in B^{+}}(a_{\mathbf{j}})^{2}$ (12)
Analogously for cross entropy with $\rho_{b}(\mathbf{x})=\sum_{\mathbf{j}\in B}b_{\mathbf{j}}\,f_{\mathbf{j}}(\mathbf{x})$:
$\int_{[0,1]^{d}}\rho(\mathbf{x})\ln(\rho_{b}(\mathbf{x}))d\mathbf{x}\approx-\sum_{\mathbf{j}\in B^{+}}a_{\mathbf{j}}b_{\mathbf{j}}$
For joint distribution of $(X,Y)$ variables (can be multivariate), denoting $B_{X},B_{Y}$ as bases used for individual variables, such approximate formula for joint entropy becomes:
$H(X,Y)=-\sum_{(\mathbf{j}_{x},\mathbf{j}_{y})\in B_{X}\times B_{Y}\backslash\{\mathbf{0}\}}\left(a_{(\mathbf{j}_{x},\mathbf{j}_{y})}\right)^{2}$ (13)
For $H(X|Y)$ conditional entropy we subtract $H(Y)$, corresponding to $\{\mathbf{0}\}\times B_{Y}^{+}$ summation indexes, leaving $B_{X}^{+}\times B_{Y}$:
$H(X|Y)=H(X,Y)-H(Y)\approx-\sum_{\mathbf{j}_{x}\in B_{X}^{+}}\sum_{\mathbf{j}_{y}\in B_{Y}}\left(a_{(\mathbf{j}_{x},\mathbf{j}_{y})}\right)^{2}$
Finally for mutual information we remove first row and column:
$I(X;Y)=H(X)+H(Y)-H(X,Y)\approx\sum_{\mathbf{j}_{x}\in B_{X}^{+}}\ \sum_{\mathbf{j}_{y}\in B_{X}^{+}}\left(a_{(\mathbf{j}_{x},\mathbf{j}_{y})}\right)^{2}$ (14)
hence, as in Fig. 7, this $\ln(1+a)\approx a$ approximation allows to evaluate mutual information by just summing squared nontrivial coefficients: mixed moments between the two variables.
However, as discussed in [24], square of estimated mean is larger than square of mean by variance divided by sample size $n$, suggesting corrected mutual information estimator:
$I_{c}(\hat{X};\hat{Y})=\sum_{\mathbf{j}\in B_{X}^{+}}\ \sum_{\mathbf{k}\in B_{Y}^{+}}\left(\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))^{2}-\frac{\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))}{n}\right)$ (15)
Figure 11: While HSIC e.g. in information bottleneck training [15] usually uses Gaussian kernel, proposed HCR approach can be also viewed as kernel method - we can see better localized in diagonal of $0,1]^{d}$ hypercube of normalized variables. While HSIC has infinite number of Gaussian centers, in HCR we use finite basis - allowing to avoid multiplication of $n\times n$ matrices for size $n$ data sample of $O(n^{3})$ or currently the lowest $O(n^{2.37})$, to linear $O(n|B|)$ for $B$ chosen basis.
Figure 12: Independence test HSIC vs HCR comparison from [24] - we independently generate 2 samples from bimodal distribution and introduce dependence by rotating it 0, 1, 2, 3, 4, 5 degrees (top). In HCR we model their joint distribution as polynomial - there are shown such density models and their $4\times 4=16$ coefficients for $m=4$. To distinguish signal from noise, they can be normalized to $N(0,1)$ assuming independence hypothesis - we can use normality test for such 16 coefficients. The most promising turned out testing the lowest and the highest, comparing them with distributions of these extreme values for 16 $N(0,1)$ variables. It turned out much more sensitive to dependencies than HSIC, also much less expensive, and additionally providing description of found dependencies.
III-B Kernel formulation like HSIC
Let us now go toward HSIC [25] kernel formulations, for simplicity without the above correction. Denoting $\{\mathbf{x}^{i}\}_{i=1..n},\{\mathbf{y}^{i}\}_{i=1..n}$ as the current batch from size $n$ data sample, denote
$\bar{X}=\frac{1}{\sqrt{n}}(f_{\mathbf{j}}(\mathbf{x}^{i}))_{i=1..n,j\in B^{+}_{X}}\qquad\bar{Y}=\frac{1}{\sqrt{n}}(f_{\mathbf{j}}(\mathbf{y}^{i}))_{i=1..n,j\in B^{+}_{Y}}$ (16)
[…]
Having above $\bar{X},\bar{Y}$ matrices of features, we can directly MSE estimate (2) parameters for all $B_{X}^{+}\times B_{Y}^{+}$ pairs forming $\theta_{XY}$ matrix. Using $\|A\|_{F}^{2}=\sum_{ij}(A_{ij})^{2}=\textrm{Tr}(AA^{T})$ and $\textrm{Tr}(AA^{\prime})=\textrm{Tr}(A^{\prime}A)$, mutual information (14) becomes:
$\theta_{XY}:=(a_{(\mathbf{j}_{x},\mathbf{j}_{y})}:\mathbf{j}_{x}\in B_{X}^{+},\mathbf{j}_{y}\in B_{Y}^{+})=\bar{X}^{T}\bar{Y}$
$\qquad I(X;Y)\approx\|\bar{X}^{T}\bar{Y}\|^{2}_{F}=\textrm{Tr}\left(\bar{X}^{T}\bar{Y}(\bar{X}^{T}\bar{Y})^{T}\right)=\textrm{Tr}(C_{X}C_{Y})$
$\textrm{for }n\times n\textrm{ {kernels}:}\quad C_{X}=\bar{X}\,\bar{X}^{T}\qquad C_{Y}=\bar{Y}\,\bar{Y}^{T}$ (17)
containing scalar products of points in batch as vectors of features. If subtracting the means earlier, they resemble covariance matrices of the used features - this subtraction might be worth including, but generally the means should be close to zero. Also, instead of covariance matrix, this is rather $n\times n$ similarity matrix inside the size $n$ batch using the chosen features.
Figure 11 shows some their values and comparison with Gaussian kernel popular in HSIC, then Fig. 12 shows their simple benchmark for the original application: independence test, where HCR turned out not only more sensitive to dependencies, and actually providing their descriptions, but is also much cheaper - avoiding costly multiplication of $n\times n$ kernels by directly using $I(X;Y)\approx\|\bar{X}^{T}\bar{Y}\|^{2}_{F}$ of $O(n|B_{X}^{+}||B_{Y}^{+}|)$ cost of mainly estimation.
III-C Gradient descent optimization of information bottleneck
The previous version of this article contains analytical solution for linear case, but in practice such training rather requires gradient as in HSIC articles ([14, 15]) we will focus on.
Denote $X\xrightarrow{\theta_{XT}}T\xrightarrow{\theta_{TY}}Y$ as the values in 3 consecutive layers: of $\times n_{X}, n×n$$\times n_{T}$, $n\times n_{Y}$ dimensions for $n$ size of dataset and $n_{X},n_{T},n_{Y}$ numbers of neurons. The optimized $\theta$ coefficients are linear dependencies between moments. For simplicity we assume being fully connected, in other case we can consider $X\to T\to Y$ triples e.g. for individual neurons in $T$, with its connections in $X$ and $Y$.
For nonlinearities we find values in the basis:
$n\times n_{T}m\ \textrm{matrix:}\quad\bar{T}_{m}=\frac{1}{\sqrt{n}}(f_{p}(t^{i}_{j}))_{i=1..n,\ (j,p)\in(1..n_{T})\times(1..m)}$
and the same for $\bar{X}_{m},\bar{Y}_{m}$. Generally we can use different basis toward left and right: $\bar{T}_{I}$ for input and $\bar{T}_{O}$ for output for some natural numbers $I,O\geq 1$, which generally can vary between layers and neurons. It means $O$ moments of deeper/earlier layer are used to predict $I$ moments of further layer, or the opposite if reversing propagation direction.
For KAN-like propagation of expected values we can choose $I=1$, or higher for propagation of probability distributions, e.g. $I=O$ being 2 to include variances, 3 skewness, 4 kurtosis. While KAN here is unidirectional ”multiple moments $\to$ expected value” propagation, coming from HCRNN we can still reverse its direction - directly using ”expected value $\to$ multiple moments”.
We would like to optimize $\bar{T}$ content of intermediate layer:
$\bar{X}_{O}\xrightarrow{\theta_{XT}=\bar{X}_{O}^{T}\bar{T}_{I}}\left(\bar{T}_{I}\xleftarrow{f_{1..I}}\bar{T}\xrightarrow{f_{1..O}}\bar{T}_{O}\right)\xrightarrow{\theta_{TY}=\bar{T}_{O}^{T}\bar{Y}_{I}}\bar{Y}_{I}$ (18)
where $\theta_{XT}=\bar{X}_{O}^{T}\bar{T}_{I}$ and $\theta_{TY}=\bar{T}_{O}^{T}\bar{Y}_{I}$ are directly MSE estimated parameters.
The chosen basis leads to mutual information approximation:
$I(X;T)\approx\textrm{Tr}(C_{X_{O}}C_{T_{I}})\quad\quad I(T;Y)\approx\textrm{Tr}(C_{T_{O}}C_{Y_{I}})$ (19)
which can be used e.g. for information bottleneck formula:
$\inf_{(t_{j}^{i}):i=1..n,\,j:1..n_{T}}(\textrm{Tr}(C_{X_{O}}C_{T_{I}})-\beta\textrm{Tr}(C_{T_{O}}C_{Y_{I}}))$ (20)
Let us find formula for this gradient starting with $C_{T},C_{T_{m}}$, for $f_{p}$ the considered orthonormal basis, $p=1,\ldots,m$:
$\frac{\partial(C_{T})_{kl}}{\partial t_{j}^{i}}=\frac{\partial(\bar{T}\bar{T}^{T})_{kl}}{\partial t_{j}^{i}}=\frac{\partial\sum_{a}t^{k}_{a}t^{l}_{a}}{n\,\partial t_{j}^{i}}=\frac{\delta_{ik}t_{j}^{l}+t_{j}^{k}\delta_{il}}{n}$
$\frac{\partial(C_{T_{m}})_{kl}}{\partial t_{j}^{i}}=\frac{\partial\sum_{ap}f_{p}(t^{k}_{a})f_{p}(t^{l}_{a})}{n\,\partial t_{j}^{i}}=\frac{\partial\sum_{p=1}^{m}f_{p}(t^{k}_{j})f_{p}(t^{l}_{j})}{n\,\partial t_{j}^{i}}=$
$\frac{\sum_{p=1}^{m}\delta_{ik}f^{\prime}_{p}(t_{j}^{i})f_{p}(t_{j}^{l})+f_{p}(t_{j}^{k})\delta_{il}f^{\prime}_{p}(t_{j}^{i})}{n}$
Mutual information e.g. between $X$ and $T$ is approximated as $\textrm{Tr}(C_{X}C_{T})$ with derivative (using symmetry of $C_{X}=(C_{X})^{T}$):
$\frac{\partial\textrm{Tr}(C_{X}C_{T})}{\partial t_{j}^{i}}=\frac{\partial\sum_{kl}C_{Xkl}C_{Tkl}}{\partial t_{j}^{i}}=\sum_{kl}C_{Xkl}\frac{\delta_{ik}t_{j}^{l}+t_{j}^{k}\delta_{il}}{n}=$
$\frac{1}{n}\sum_{l}C_{Xil}t_{j}^{l}+\frac{1}{n}\sum_{k}C_{Xki}t_{j}^{k}=\frac{2}{n}\sum_{k}C_{Xik}t_{j}^{k}=\frac{2}{\sqrt{n}}(C_{X}\bar{T})_{ij}$
$\left(\frac{\partial I(X;T)}{\partial t_{j}^{i}}\approx\right)\quad\frac{\partial\textrm{Tr}\,(C_{X}C_{T_{m}})}{\partial t_{j}^{i}}=\sum_{kl}C_{Xkl}\frac{\partial(C_{T_{m}})_{kl}}{\partial t_{j}^{i}}=$
$\frac{2}{n}\sum_{kp}C_{Xik}\,f_{p}(t_{j}^{k})f^{\prime}_{p}(t_{j}^{i})=\frac{2}{\sqrt{n}}\sum_{p=1}^{m}(C_{X}\tilde{T}_{p})_{ij}\,f^{\prime}_{p}(t^{i}_{j})$
for $n\times n_{T}$ matrix $\tilde{T}_{p}=\frac{1}{\sqrt{n}}(f_{p}(x^{i}_{j}))_{i=1..n,j=1..n_{T}}$ containing single power, while $n\times mn_{T}$ matrix $\bar{T}_{m}$ contains all $\tilde{T}_{p=1..m}$. Denoting $\tilde{T}^{\prime}_{p}=\frac{1}{\sqrt{n}}(f^{\prime}_{p}(t^{i}_{j}))_{i=1..n,j=1..n_{T}}$, the above become coordinatewise matrix multiplication.
The final gradient of optimized (20) information bottleneck function, also extending $C_{X_{O}}=\bar{X}_{O}\bar{X}_{O}^{T},C_{Y_{I}}=\bar{Y}_{I}\bar{Y}_{I}^{T}$ (e.g. to reduce cost by multiplying by $\tilde{T}_{p}$ first), becomes:
$\frac{\partial\,(\textrm{Tr}(C_{X_{O}}C_{T_{I}})-\beta\textrm{Tr}(C_{T_{O}}C_{Y_{I}}))}{\partial t_{j}^{i}}=$ (21)
$\frac{2}{\sqrt{n}}\left(\sum_{p=1}^{I}\left(\bar{X}_{O}\,\bar{X}_{O}^{T}\tilde{T}_{p}\right)_{ij}f^{\prime}_{p}(t^{i}_{j})-\beta\sum_{p=1}^{O}\left(\bar{Y}_{I}\,\bar{Y}_{I}^{T}\tilde{T}_{p}\right)_{ij}f^{\prime}_{p}(t^{i}_{j})\right)$
where generally there can be used various $I,O$ basis size for different layers, neurons. Above multiplications could use e.g. $(\bar{X}_{O}\,\bar{X}_{O}^{T})\tilde{T}_{p}$ or $\bar{X}_{O}\,(\bar{X}_{O}^{T}\tilde{T}_{p})$ bracketing - while the former uses $n\times n$ kernel, the latter is less expensive over basis - allowed for HCR, but not HSIC.
For practical trening we need to split dataset into batches applied multiple times (epochs) in some sequence, e.g. starting with random initial values of weights/intermediate layers. There is a large freedom to choose training details, where we can apply two philosophies:
III-C1 Modification of weights based on gradients
We can use the formulas for weight estimation from given batch: $\theta_{XT}=\bar{X}_{O}^{T}\bar{T}_{I}$, $\theta_{TY}=\bar{T}_{O}^{T}\bar{Y}_{I}$, for example to update the actually used weights $\bar{\theta}$, e.g. with exponential moving average: $\bar{\theta}+=\eta(\theta-\bar{\theta})$. Alternatively we can use the found $g_{ij}=\partial_{t_{ij}}$ gradient, and apply it to modify parameters using $\bar{T}\to\bar{T}-\alpha g$.
III-C2 Modifying content of intermediate layers
We could use this gradient to directly modify contents of intermediate layers corresponding to datapoints. However, as such normalized variables have to stay in $[0,1]$ range, direct use of gradient descent method could easily leave this range. To prevent that, we can for example stretch the variable to $(-\infty,\infty)$ for gradient descent step, e.g. with $s=\textrm{CDF}^{-1}(t),t=\textrm{CDF}(s)$ using CDF of e.g. Gaussian distribution, transforming these hidden variables to this distribution. Derivative over the stretched variable $s$ needs just additional multiplication by its PDF: $\partial L(\textrm{CDF}(s))/\partial s=L^{\prime}(t)\,\textrm{CDF}^{\prime}(s)$.
IV Conclusions and further work
Neurons with joint distribution models seem powerful agnostic improvement for currently popular guessed parametrizations like MLP or KAN, and are practically accessible with HCR, up to omnidirectional neurons like in Fig. 13 - allowing to freely choose inference directions, propagate both values and probability distributions, with clear coefficient interpretations.
As in Fig. 3, 1, BNNs are qualitatively superior than current ANNs, also have multidirectional propagation, including probabilities, and need different training than standard backpropagation. Proposed new ANNs allow to catch up with such low level behavior still looking biologically plausible e.g. KAN-like, hopefully allowing to also get closer for high level behavior, maybe recreating mathematics hidden in BNN behavior.
However, mastering such new neural network architecture will require a lot of work, planned also for future versions of this article. Here are some basic research directions:
  • Search for practical applications, from replacement of standard ANN, for multidirectional inference e.g. in Bayes-like scenarios, as neural networks propagating probability distributions, up to exploration of similarity/replacement for biological neurons.
  • Practical implementation, optimization especially of training and update, basis optimization and selection techniques, exploration of tensor decomposition approach.
  • Working on probability distributions makes it natural for information theoretic approaches like information bottleneck [5] optimization of intermediate layers, also hopefully leading to better understanding e.g. of information propagation during learning/inference, information held by intermediate layers, etc.
  • Adding time dependence like model update, also for similarity with biological neurons, e.g. long term potentiation, connection to various periodic processes/clocks.
  • While the discussed neurons containing joint distribution models seem very powerful and flexible, directly working in high dimensions they have various issues - suggesting to directly predict conditional distributions instead with HCR parametrization ([9, 10, 12, 13]), what might be also worth included in neural network, e.g. as a part of the training process - to be decomposed into single neurons.
  • Modification of softmax e.g. embedding in transformer approach from single parameter to multiple HCR moments representing probability distribution of embedded parameter.
Figure 13: Omnidirectional HCR neuron proposed in [8] - getting any subset $S$ of connections as input, it can update model of joint distribution inside $S$: for $a_{\mathbf{j}}$ coefficients positive only in this subset $\left(\{i:j_{i}\geq 1\}\subset S\right)$, and predict/propagate to output as the remaining connections e.g expected values for these inputs, for example accumulated up to some threshold, including sign for excitatory/inhibitory.